An astronomical telescope of ten fold angular magnification has a length of the 44 cm. The focal length of the objective is. 
A. 4cm
B.40cm
C.44cm
D . 440 cm

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Shuchi
 
Given: Length of astronomical telescope (f0+ fe) = 44 cm and magnification

(f_0/f_e )=10.

From the given magnification, we find that f0 = 10 fe.

Therefore, 10 fe + fe = 44 or 11 fe = 44 or fe= 4.

And focal length of the objective (f0) = 44 - fe = 44 - 4 = 40 cm.

 

Regards

Arun (askIITians forum expert)